We tried to solve this type of problem.....thought we were correct, however the teacher said otherwise. Of course she said we could "re-do" the incorrect problems, but she gave no further explanation. Use prime factors to simplify each expression.
4 4/15 + 2 4/39=

Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

4 4/15 + 2 4/39=
---------------------------
(15*4+4)/15 + (39*2+4)/39
-------------------------------
(64/15) + (82/39)
---
Use prime factor form to find the least common denominator:
----
15 = 3*5
39 = 3*13
------
The lcm is 3*5*13 = 195
--------------------------
Rewrite each factor with the lcm as its denominator:
(64*13/lcm) + (82*5/lcm)
------------------------------
Combine the numerators over the lcm:
(832 + 410)/lcm
------
1242/(195)
=================
Cheers,
Stan H.

Answer by checkley77(12844)

4 4/15 + 2 4/39
(4*15+4)/15+(2*39+4)/39 chNge to improper fractions.
(60+4)/15+(78+4)/39
64/15+82/39 find common denominator & combine.
(64*39+82*15)/15*39
(2,496+1,230)/585
3,726/585 reduce divide by 9/9.
414/65 ans.